OpenStudy (anonymous):

Copper metal (Cu) reacts with silver nitrate (AgNO3) in aqueous solution to form Ag and Cu(NO3)2. An excess of AgNO3 is present. The balanced chemical equation is shown below. Cu + 2AgNO3 ---> Cu(NO3)2 + 2Ag The molar mass of Cu is 63.5 g/mol. The molar mass of Ag is 107.9 g/mol. What mass, in grams, of Ag is produced from reaction of 31.75 g of Cu? 26.95 107.9 215.91 431.82

OpenStudy (aaronq):

So were told that Cu is the limiting reactant (because silver nitrate is in excess). The next step is to convert the mass (grams) into moles by using the molar mass. \(\sf moles =\dfrac{mass}{Molar ~mass}\) Next, we use the stoichiometric coefficients (numbers infront) (and moles of what were interested in) in a ratio: \(\sf \dfrac{moles~of~Cu}{Cu's~coefficient}=\dfrac{moles~of~Ag}{Ag's~coefficient}\) Solve for moles of Ag. Next, convert the moles of Ag to mass. Again, use the formula: \(\sf moles =\dfrac{mass}{Molar ~mass}\)

OpenStudy (anonymous):

1m of ag?

OpenStudy (aaronq):

i dont know, have you done the calculations?

OpenStudy (anonymous):

I got .5m of Cu.

OpenStudy (aaronq):

whats m?

OpenStudy (anonymous):

Mole*

OpenStudy (aaronq):

oh okay, write "mol", "m" means molality How many moles of Ag did you get?

OpenStudy (anonymous):

1

OpenStudy (aaronq):

perfect. now you converted to grams?

OpenStudy (anonymous):

1 mol Ag = 107.9g

OpenStudy (aaronq):

thats it! good job